# Beyond Rate Coding: Surrogate Gradients Enable Spike Timing Learning in Spiking Neural Networks

### v1

@@ Abstract
  • We investigate the extent to which Spiking Neural Networks (SNNs) trained with Surrogate Gradient Descent (Surrogate GD), with and without delay learning, can learn from precise spike timing beyond firing rates. We first design synthetic tasks isolating intra-neuron inter-spike intervals and cross-neuron synchrony under matched spike counts. On more complex spike-based speech recognition datasets (Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC), we construct variants where spike count information is eliminated and only timing information remains, and show that Surrogate GD-trained SNNs are able to perform significantly above chance whereas purely rate-based models perform at chance level. We further evaluate robustness under biologically inspired perturbations – including Gaussian jitter per spike or per-neuron, and spike deletion – revealing consistent but perturbation-specific degradation. Networks show a sharp performance drop when spike sequences are reversed in time, with a larger drop in performance from SNNs trained with delays, indicating that these networks are more human-like in terms of behaviour. To facilitate further studies of temporal coding, we have released our modified SHD and SSC datasets.
@@ 1 Introduction
Spiking Neural Networks (SNNs) are biologically inspired models that compute via discrete, sparse spikes, rather than continuous activations (Maass, 1997). This event-driven framework not only captures rich temporal patterns (such as inter-spike intervals (ISI), cross-neuron synchrony, causal spike sequences (Bohte, 2004)) but also powers energy-efficient neuromorphic hardware (Painkras et al., 2013; Akopyan et al., 2015; Davies et al., 2018; Orchard et al., 2021; Su et al., 2024).
@@
Despite their promise, a persistent challenge remains: how to convert precise spike timing into effective learning signals. Surrogate Gradient Descent (Surrogate GD) provides a practical solution by replacing the non-differentiable spike with a smooth surrogate, thereby allowing backpropagation through time (Neftci et al., 2019). Subsequent work has improved its efficiency and adaptiveness by enforcing sparser gradients (Perez-Nieves and Goodman, 2021) and by developing more adaptive surrogate functions (Wang et al., 2023). Surrogate GD–trained networks have matched or outperformed rate-based RNNs on several neuromorphic benchmarks (Shrestha and Orchard, 2018; Dampfhoffer and Mesquida, 2024; Hammouamri et al., 2024; Sun et al., 2025), but it remains unclear whether these successes stem from genuine temporal coding or from reliance on coarse firing-rate features.
@@
To address this ambiguity, researchers have injected various temporal inductive biases: heterogeneous time constants (Perez-Nieves et al., 2021; Habashy et al., 2024); learnable axonal delays (Shrestha et al., 2022; Sun et al., 2022; Patiño-Saucedo et al., 2023); synaptic delays (Zhang et al., 2020; Yu et al., 2022; Sun et al., 2024; Hammouamri et al., 2024); and dendritic delay models (D’agostino et al., 2024; Zheng et al., 2024). Yet even with these enhancements, Surrogate GD alone may lack the bias needed to capture precise timing.
@@
This uncertainty leads to a central question: Can Surrogate GD-based SNNs learn and depend on spike timing codes beyond rate statistics? To address this, we investigate this question through a multi-stage evaluation. First, we construct synthetic benchmark tasks that isolate the contribution of ISI and spike synchrony under controlled spike counts, building upon prior work that highlights the ability of SNNs to leverage temporal coding for classification (Bohte et al., 2002; Mirsadeghi et al., 2021). Then, we evaluate models on datasets with real-world temporal complexity (SHD and SSC) (Cramer et al., 2020), introducing spike count-normalized variants to remove rate confounds. Finally, we assess robustness under biologically inspired perturbations—Gaussian jitter applied per-spike and per-neuron (Amarasingham et al., 2012; Agmon, 2012), spike deletions (Mainen and Sejnowski, 1995; Chen et al., 2022), and time reversal.
@@
Our findings show that Surrogate GD-trained SNNs can extract fine-grained temporal structure even without spike rate information, including intra-neuron ISIs and cross-neuron synchrony. Delay learning further enhances the model’s ability to exploit such timing features. Across perturbations—per-spike and per-neuron Gaussian jitter, and per-spike deletion—we observe distinct degradation patterns that reveal both vulnerabilities and robustness in the model’s temporal encoding. Notably, under time reversal, Surrogate GD-trained SNNs not only leverage reversal-invariant features like ISI and coincidence, but also capture temporal dependencies in spike timing, indicating a layered temporal representation.
@@ 2 Surrogate Gradient Descent Enables Learning of Diverse Spike-Timing Codes
Throughout this section we analyse two synthetic benchmarks that isolate specific temporal features—intra-neuron inter-spike intervals (ISI) and inter-neuron coincidence. Forward propagation uses a Heaviside spike function, while back-propagation employs the smooth surrogate
$\frac{\partial H(x)}{\partial x}\approx\frac{1}{(\alpha\cdot|x|+1)^{2}},\quad\alpha=100,$ (1)
@@
following the general form proposed by (Neftci et al., 2019). Unless otherwise noted, it is used consistently across all SGD models in this work.
Figure 1: ISI-based dataset. (a) Class decision boundary in ISI-rate space. Horizontal rate-only cutoff (gray) achieves only partial separation. (b) Example spike trains generated using fixed ISI pairs (top) and their randomized counterparts with $f=0.5$ (bottom). (c) Test accuracy under increasing spike timing perturbation $f$ (fraction of spike times randomly redrawn). Accuracy degrades toward the spike count baseline as $f\to 1$.
2.1 Surrogate GD Successfully Learns ISI-Based Temporal Codes
@@
This experiment evaluates whether SNNs trained with surrogate gradients can learn from fine-grained inter-spike interval (ISI) structure. Each sample contains $N=10$ input neurons, each generating spike pairs with a fixed class-specific ISI $\delta$ and firing rate $r$. Spikes are distributed within a $T=1000$ ms window and placed at non-overlapping positions. The number of spikes and ISI steps are computed by:
$\text{Number of spike pairs}=\left\lfloor r\cdot\frac{T}{1000}\right\rfloor,\quad\text{ISI steps}=\max\left(1,\left\lfloor\delta\right\rfloor\right).$ (2)
The model consists of an input layer, a spiking hidden layer with 100 neurons and a shared learnable membrane time constant $\tau$. The model is non-recurrent, and classification is based on output membrane potentials at the final time step using cross-entropy loss.
Random Replacement. To simulate timing degradation, a fraction $f\in[0,1]$ of spikes in each neuron’s train are removed and replaced by new spikes sampled uniformly over $[0,T)$:
$S_{i}^{\prime}=(S_{i}\setminus\mathcal{R}_{i})\cup\mathcal{U}_{i},\quad|\mathcal{R}_{i}|=|\mathcal{U}_{i}|=\left\lfloor f\cdot|S_{i}|\right\rfloor.$ (3)
As shown in fig. 1(c), the model achieves approximately 95% accuracy under clean input conditions ($f=0$), successfully solving the ISI classification task. This confirms that Surrogate Gradient Descent enables the SNN to extract information from inter-spike intervals—one of the fundamental forms of spike timing codes.
@@
As the perturbation factor $f$ increases, spike timing information becomes increasingly disrupted, leading to degraded model performance. For $f>0.5$, most ISIs are broken—as each interval requires a pair of spikes—effectively eliminating temporal cues. The model then relies primarily on firing rate, with accuracy converging to approximately 75%. This is above chance level (50%) but reflects the limit of rate-based discrimination.
2.2 Surrogate GD Successfully Learns Coincidence-Based Temporal Codes
This experiment evaluates whether SNNs trained with surrogate gradient descent can classify inputs based on group-level synchrony patterns. Each sample consists of $N=60$ input neurons evenly divided into three groups. Spikes are generated in non-overlapping windows of size $w=10$ within a total duration $T$, and are restricted to a central sub-window of 5 steps to introduce mild temporal jitter while maintaining synchrony.
The class identity is defined by which two of the three groups fire synchronously while the third inverts. For example, Class A corresponds to synchrony between Groups 1 and 2, with Group 3 in OFF state. To eliminate rate-based cues, ON/OFF group activity toggles randomly in each window, and total spike count is balanced across samples.
Formally, for each window $w_{i}$, let $\mathcal{T}_{i}=\{t_{i}+2,\dots,t_{i}+6\}$ be the valid spike time slots. For each group $g$ and class $c$:
$k_{i}^{(g)}\sim P_{c}(k;\mu_{\text{on/off}}),\quad\text{activate }k_{i}^{(g)}\text{ neurons in group }g\text{ at times }t\in\mathcal{T}_{i}.$ (4)
To control task difficulty, we interpolate between ON/OFF spike count distributions via a synchrony overlap factor $\lambda\in[0,1]$:
$\mu_{\text{on}}=(1-\lambda)\cdot\mu^{(0)}_{\text{on}}+\lambda\cdot\mu_{\text{avg}},\quad\mu_{\text{off}}=(1-\lambda)\cdot\mu^{(0)}_{\text{off}}+\lambda\cdot\mu_{\text{avg}},$ (5)
@@ 2.2 Surrogate GD Successfully Learns Coincidence-Based Temporal Codes
where $\mu_{\text{on}}^{(0)}=12$, $\mu_{\text{off}}^{(0)}=2$, and $\mu_{\text{avg}}=5$. As $\lambda$ increases, the ON and OFF spike distributions converge, reducing synchrony discriminability (see fig. 2(a)).
@@
Figure 2: Coincidence-based dataset. (a) Synchronous spike count distributions for ON/OFF classes at $\lambda=0$ and $\lambda=1$. As $\lambda$ increases, ON and OFF distributions converge. (b) Spike raster plots for each class under low ($\lambda=0$) and moderate ($\lambda=0.5$) synchrony overlap. Red lines mark group boundaries. Light gray and dark gray bands indicate ON and OFF neuron groups, respectively. (c) Test accuracy of the Surrogate GD trained SNN across varying levels of synchrony overlap $\lambda$.
We employ a compact architecture with a 60-neuron input layer, a 3-neuron spiking hidden layer (one per group), and a 3-neuron non-spiking output layer. This minimal hidden size allows direct interpretation of how synchrony patterns map to class labels. Classification is based on the final membrane potentials of the output layer using cross-entropy loss.
As shown in fig. 2(c), the model achieves near-perfect accuracy at $\lambda=0$, indicating successful learning of group-level spike synchrony. Accuracy declines as $\lambda$ increases and synchrony degrades, converging toward the rate-based baseline of 33.3%.
3 Validation of Spike Timing Learning on SHD/SSC Datasets
Motivated by our synthetic-task results showing that Surrogate GD-trained SNNs can leverage timing information, we evaluate them on more complex speech recognition datasets—SHD and SSC—which exhibit pronounced spike-count biases that allow MLPs to achieve nontrivial accuracy. To isolate genuine timing-based learning, we therefore construct two timing-normalized variants (see fig. 3) that remove rate information while preserving each sample’s temporal structure.

### v2

@@ Abstract
  • The surrogate gradient descent algorithm enabled spiking neural networks to be trained to carry out challenging sensory processing tasks, an important step in understanding how spikes contribute to neural computations. However, it is unclear the extent to which these algorithms fully explore the space of possible spiking solutions to problems. We investigated whether spiking networks trained with surrogate gradient descent can learn to make use of information that is only encoded in the timing and not the rate of spikes. We constructed synthetic datasets with a range of types of spike timing information (interspike intervals, spatio-temporal spike patterns or polychrony, and coincidence codes). We find that surrogate gradient descent training can extract all of these types of information. In more realistic speech-based datasets, both timing and rate information is present. We therefore constructed variants of these datasets in which all rate information is removed, and find that surrogate gradient descent can still perform well. We tested all networks both with and without trainable axonal delays. We find that delays can give a significant increase in performance, particularly for more challenging tasks. To determine what types of spike timing information are being used by the networks trained on the speech-based tasks, we test these networks on time-reversed spikes which perturb spatio-temporal spike patterns but leave interspike intervals and coincidence information unchanged. We find that when axonal delays are not used, networks perform well under time reversal, whereas networks trained with delays perform poorly. This suggests that spiking neural networks with delays are better able to exploit temporal structure. Since humans also perform poorly with time-reversed speech, it suggests that these networks may be better suited as neural models. To facilitate further studies of temporal coding, we have released our modified speech-based datasets.
@@ 1 Introduction
Spiking Neural Networks (SNNs) are biologically inspired models that compute via discrete, sparse spikes, rather than continuous activations (Maass, 1997). This event‑driven framework not only captures rich temporal patterns (such as inter‑spike intervals and cross‑neuron synchrony) but also powers energy‑efficient neuromorphic hardware (Painkras et al., 2013; Akopyan et al., 2015; Davies et al., 2018; Orchard et al., 2021; Su et al., 2024).
@@
Despite their promise, a persistent challenge remains: how to convert precise spike timing into effective learning signals. Surrogate Gradient Descent (Surrogate GD) provides a practical solution by replacing the non‑differentiable spike with a smooth surrogate, thereby allowing backpropagation through time (Neftci et al., 2019). Subsequent work has improved its efficiency and adaptiveness by enforcing sparser gradients (Perez-Nieves and Goodman, 2021) and by developing more adaptive surrogate functions (Wang et al., 2023). Surrogate GD–trained networks have matched or outperformed rate‑based recurrent neural networks on several neuromorphic benchmarks (Shrestha and Orchard, 2018; Sun et al., 2025a; Dampfhoffer and Mesquida, 2024; Hammouamri et al., 2024; Sun et al., 2025b), but it remains unclear whether these successes stem from genuine temporal coding or from reliance on coarse firing‑rate features.
@@
To address this ambiguity, researchers have injected various temporal inductive biases: heterogeneous time constants (Fang et al., 2021; Perez-Nieves et al., 2021; Habashy et al., 2024); learnable axonal delays (Shrestha et al., 2022; Sun et al., 2022; Patiño-Saucedo et al., 2023); synaptic delays (Zhang et al., 2020; Yu et al., 2022; Sun et al., 2024; Hammouamri et al., 2024); and dendritic delay models (D’agostino et al., 2024; Zheng et al., 2024). Yet even with these enhancements, there has been no systematic test of whether or not Surrogate GD finds networks that make use of precise timing.
@@
This uncertainty leads to a central question: can Surrogate GD-trained SNNs learn and depend on spike timing codes beyond rate statistics? This question is important because if they were just using rate-based information, then (a) they are not making use of all the information that is available, and are therefore less efficient than they could be, (b) it limits their effectiveness as a model of spike-based computation in the brain, and (c) it raises the question of whether or not there is an advantage of using such expensive training methods compared to simply approximating rate-based networks. To address this central question, we used both synthetic datasets as well as more complicated real-world benchmarks. First, we construct synthetic benchmark tasks that isolate the contribution of inter-spike interval (ISI), coincidence, and cross-channel inter-spike interval (CCISI, a type of spatio-temporal spike pattern or polychrony) under controlled spike counts, building upon prior work that highlights the ability of SNNs to leverage temporal coding for classification (Bohte et al., 2002; Mirsadeghi et al., 2021). Then, we evaluate models on datasets with real-world temporal complexity (Cramer et al., 2020), introducing spike count-normalized variants to remove rate confounds. We assess robustness under biologically inspired perturbations: Gaussian jitter applied per-spike and per-neuron (Amarasingham et al., 2012; Agmon, 2012), and spike deletions (Mainen and Sejnowski, 1995; Chen et al., 2022). Finally, we analyse the effect of time reversal that selectively disrupts spike patterns while leaving coincidences and ISIs unchanged.
@@
Our findings show that Surrogate GD-trained SNNs can extract fine-grained temporal structure even without spike rate information, including intra-neuron ISIs, cross-neuron synchrony or coincidence, and cross-channel causal ISI. Delay learning further enhances the model’s ability to exploit such timing features, especially for long-range causal gaps that exceed the membrane integration window controlled by the time constant. Across perturbations we observe distinct degradation patterns that reveal both vulnerabilities and robustness in the model’s temporal encoding. Notably, under time reversal, Surrogate GD-trained SNNs not only leverage reversal-invariant features like ISI and coincidence, but also capture causal dependencies in spike timing (as in CCISI), indicating a layered temporal representation.
@@ 2 Surrogate Gradient Descent Enables Learning of Diverse Spike-Timing Codes
Throughout this section we analyse three synthetic benchmarks that isolate specific temporal features—intra-neuron inter-spike intervals (ISI), inter-neuron coincidence, and cross-channel inter-spike intervals (CCISI).
2.1 Models
The model is a feedforward spiking neural network, consisting of an input layer of $N$ neurons, a fully connected spiking hidden layer (with 100 neurons for the ISI and CCISI tasks, see sections˜2.2 and 2.3, and 3 neurons for the coincidence task, see section˜2.4), and a fully connected spiking output layer whose size equals the number of classes. The membrane time constant $\tau$ is shared between all neurons, and learned during training. We compare two architectures: (i) synaptic weights and $\tau$ are learnable, hereafter referred to as Learnable Tau; and (ii) the same model augmented with learnable axonal delays, referred to as Tau + Delay.
Training is implemented in the SLAYER framework (Shrestha and Orchard, 2018). In this framework, during forward propagation, the spike threshold uses a Heaviside function $H(x)$, while during back-propagation it employs the smooth surrogate
$\frac{\mathrm{d}H(x)}{\mathrm{d}x}\mathrel{\stackrel{{\scriptstyle\triangle}}{{=}}}\frac{1}{(\alpha\cdot|x|+1)^{2}},\quad\alpha=100,$ (1)
@@ 2.1 Models
following the general form proposed by Neftci et al. (2019). Unless otherwise noted, this surrogate gradient is used consistently across all models.
During training, classification is supervised with a ProbSpikes cross-entropy loss on the output membrane potentials, and at inference, the predicted label is given by the index of the output neuron that fired the most spikes over the simulation time. ProbSpikes loss computes cross-entropy on normalized spike counts rather than membrane potentials: the spikes from each output neuron over the time window are converted into a probability distribution and compared with the one-hot target. This encourages the correct class to emit the largest fraction of spikes (Shrestha et al., 2022).
2.2 Surrogate GD Learns Inter-Spike Intervals
We first evaluate whether SNNs trained with surrogate gradients can learn from fine-grained inter-spike interval (ISI) structure. We do this by constructing a dataset where it is impossible to perform much above chance level without using ISIs.
@@ 2.2 Surrogate GD Learns Inter-Spike Intervals
In our dataset, each sample contains $N=10$ input neurons, with each neuron generating spike pairs with a fixed class-specific ISI $\delta$ and firing rate $r$ (fig.˜1(a)). Spikes are distributed within a window and placed at non-overlapping positions (fig.˜1(b) upper). The number of spikes is $r\cdot\tfrac{T}{1000}$. When the max interval is less than 100 ms, we use $T=1,000$ and $r\in[1,50]$, and for longer intervals we use $T=10,000$ and $r\in[0.1,1]$ to ensure that placement of all pairs is possible.
To simulate timing degradation, a fraction $f\in[0,1]$ of spikes in each neuron’s train are removed and replaced by new spikes sampled uniformly over $[0,T)$. This procedure ensures that the total number of spikes remains unchanged and that no two spikes overlap in time for the same neuron. When $f=0$ all spikes come in pairs with an unambiguous ISI. When $f=1$, the spike train is fully randomized, removing all ISI information while still preserving the original spike count.
@@
In the clean case where $f=0$ the models achieve approximately 100% accuracy (fig.˜1(d)), demonstrating that surrogate gradient descent is able to extract temporal structure from spike trains encoded in the form of interspike intervals. Models with and without learnable delays perform about equally across the whole range of perturbation levels $f$, indicating that the delay module does not provide additional benefit in this setting. As the perturbation factor $f$ increases from 0 to 1, spike timing information becomes increasingly disrupted, leading to degraded model performance, although performance remains close to perfect until $f=0.2$ (20% of spikes disrupted) and only substantially degrades for $f\geq 0.4$. For $f>0.5$ (fig.˜1(b) lower), most ISIs are broken, as each interval requires a pair of spikes. The model then relies primarily on firing rate, with accuracy converging to approximately 60%. This is above chance level (50%) but reflects the limit of rate-based discrimination.
Figure 1: ISI and CCISI-based datasets. (a) Shared feature space spanned by firing rate and ISI/CCISI interval. A full decision boundary (black) separates classes more effectively than a rate-only threshold (gray). (b) Example ISI spike trains with fixed intervals (top) and their randomized versions with $f=0.5$ (bottom). (c) Example CCISI spike trains with fixed cross-neuron delays (top) and their randomized counterparts with $f=0.5$ (bottom). (d) Test accuracy under increasing spike timing perturbation $f$ (evaluated at maximum interval of 50 ms), comparing models with a learnable membrane constant $\tau$ and those augmented with delays. (e) Accuracy across increasing maximum ISI/CCISI intervals.
2.3 Delay Module Enhances Cross-Channel ISI Code Learning
To further probe the role of delay learning, we designed a cross-channel ISI (CCISI) task in which spike timing is explicitly causal across neuron pairs. In this setup, neurons are arranged into pairs $(a,b)$, where neuron $a$ fires first and neuron $b$ follows after an interval $\delta$. This enforces a directional temporal dependency that cannot be reduced to firing rate or single-neuron ISI statistics.
Formally, each sample contains $N$ neurons, where $N$ is even, grouped into fixed disjoint pairs $(2i,2i+1)$ for $i=0,\dots,N/2-1$. For a firing rate $r$ (Hz) and total duration $T$ (ms), the number of spike pairs is $r\cdot\tfrac{2T}{1000}$, and we use the same duration $T$ and range of values of $r$ as in the previous section.
For each pair $(a,b)$, neuron $a$ emits a spike at time $t$, and neuron $b$ is constrained to fire at $t+\delta$, with the additional constraint that no two spike pairs overlap within the interval $t-\delta,\,t+\delta]$ to avoid conflicts.
In the same setup as the previous section, we get similar results for CCISI as we did for the ISI task (fig.˜1(d)). However, we find that when we increase the maximum CCISI from 50 ms to up to 500 ms strong differences emerge between these tasks and the two models (Learnable Tau and Tau + Delay). For the ISI task, both models perform similarly over this range (fig.˜1(e)). For CCISI however, the difference between the two models becomes pronounced at longer intervals.
This difference between the models with and without delays at longer intervals may be a result of the fact that the membrane time constant is more constrained in having to be long enough to integrate information across all timescales that are present, but short enough to be sensitive to differences on those timescales. By contrast, delays are not so constrained, and are therefore easier to adapt to tasks in which a wider range of timescales are present. We will return to this in section˜3.
2.4 Surrogate GD Enables Coincidence-Based Temporal Code Learning
We next turn to the question of whether or not surrogate gradient descent can learn to extract information encoded in synchrony statistics (rather than precise spike patterns).
In this task, each sample consists of $N=60$ input neurons evenly divided into three groups (2(b)). Time is divided into windows of length $w=200$ time steps, and in each window a group is either ON (high mean firing rate $\mu_{\text{on}}$) or OFF (low mean firing rate $\mu_{\text{off}}$). The class identity is defined by which two of the three groups fire with synchronous patterns (in every window, both are ON or both OFF) while the third inverts (OFF when the other two groups are ON, and vice versa). For example, Class A corresponds to synchrony between Groups 1 and 2, with Group 3 out of synchrony. To eliminate rate-based cues, ON/OFF group activity toggles randomly in each window, and so the total spike count is balanced across groups and samples, meaning that firing rate gives no information about class. Spike counts are Poisson distributed, and spikes are randomly allocated to times in the central 50% of each window. To control task difficulty, we interpolate between ON/OFF spike count distributions via a synchrony overlap factor $\lambda\in[0,1]$:
$\mu_{\text{on}}$ $=(1-\lambda)\cdot\mu^{(0)}_{\text{on}}+\lambda\cdot\mu_{\text{avg}},$ (2)
$\mu_{\text{off}}$ $=(1-\lambda)\cdot\mu^{(0)}_{\text{off}}+\lambda\cdot\mu_{\text{avg}},$
@@ 2.4 Surrogate GD Enables Coincidence-Based Temporal Code Learning
where $\mu_{\text{on}}^{(0)}=12$, $\mu_{\text{off}}^{(0)}=2$, and $\mu_{\text{avg}}=5$. As $\lambda$ increases, the ON and OFF spike distributions converge, reducing synchrony discriminability (fig.˜2(a)).
@@
Figure 2: Coincidence-based dataset. (a) Spike count distributions for ON/OFF windows at $\lambda=0$ and $\lambda=1$. As $\lambda$ increases, ON and OFF distributions converge. (b) Spike raster plots for each class under low ($\lambda=0$) and moderate ($\lambda=0.5$) synchrony overlap. Red lines mark group boundaries. Light gray and dark gray bands indicate ON and OFF neuron groups, respectively. (c) Test accuracy across varying synchrony overlap $\lambda$, comparing models with a learnable $\tau$ versus $\tau$+delay.
Again, surrogate gradient descent can learn to solve this task with near perfect accuracy in the clean case ($\lambda=0$), and degrades to chance level as $\lambda\rightarrow 1$ (fig.˜2(c)). There is little to no difference between the models (with and without delays), unsurprisingly given that time differences play no part in this task.
3 Validation of Spike Timing Learning on Speech Datasets
We have seen that surrogate gradient descent can learn to extract various forms of temporal structure (synchrony and ordered/unordered time differences within/across channels) in controlled conditions. We now test whether surrogate gradient–trained SNNs can exploit spike timing information in more realistic sensory tasks. To this end, we evaluate them on SHD and SSC, which consist of samples of single spoken words from 20/35 classes (respectively) converted to spike trains via a simulated cochlea model (Cramer et al., 2020). These datesets exhibit complex spatiotemporal structure but also strong spike-count variation that allows rate-based models to perform surprisingly well. To disentangle rate and timing contributions, we construct two timing-normalized dataset variants (fig.˜3) that progressively remove rate information while preserving each sample’s temporal structure.
3.1 Models
All experiments in this section use a two-layer feedforward SNN (fig.˜4) implemented in the SLAYER framework (Shrestha and Orchard, 2018), with 128 neurons per layer and the Spikemax loss function (Shrestha et al., 2022). We compare two variants: a baseline model trained by surrogate gradient descent (SGD), which learns synaptic weights and an enhanced version (SGD-delay) that additionally learns axonal delays (Sun et al., 2023). Predictions are obtained from the neuron with the highest spike count over the simulation window.
We also compare to the performance of a multilayer perceptron (MLP) using a vector of spike counts for each neuron as input. This serves as a proxy for the performance level attainable using only firing rates, and not spike times. The MLP comprises two fully connected layers with 256 hidden units and ReLU activation, mapping the spike count vector to a linear readout over all output classes.